There are 17 math team members, they must pick five members to go to regional how many combinations are possible? A.170 B.12,376 C.85 D.6188
Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

There are 17 math team members, they must pick five members to go to regional how many combinations are possible? A.170 B.12,376 C.85 D.6188
---
17C5 = (17*16*15*14*13)/(1*2*3*4*5) = 6188
=================
Cheers,
Stan H.

Answer by jim_thompson5910(35256)

Assuming order does not matter, there are 17 C 5 = 17!/(5!(17-5)!) = 17!/(5!12!) = (17*16*15*14*13*12!)/(5!12!) = (17*16*15*14*13)/(5*4*3*2*1) = 6188 different combinations.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim